# Label-Free Multivariate Time Series Anomaly Detection

### 17 Dec 2023 v1

@@ 1 Introduction
Density estimation is a promising approach for unsupervised anomaly detection because they do not depend on the assumption that training datasets are all normal. As presented in Fig. 1b, it detects an abnormal sample based on the widely accepted hypothesis that abnormal instances exhibit sparse densities than the normal, i.e., abnormal samples are typically in the low-density regions while normal samples are in the high-density regions [25, 26, 27]. Previous works have attempted this strategy for the prediction task and the key challenge lies in the accurate density estimation of the distribution. In [28, 29, 30], time series density is modeled as the parameterized probability distribution, while it is still challenging to model a more complex data distribution [31].
@@
Previous work GANF [32] tackles the same MTS anomaly detection task. In their design, the static directed acyclic graph (DAG) is leveraged to model intractable dependence among multiple entities, and normalizing flow is employed to estimate an overall distribution for all entities together [33, 34]. Although GANF has achieved state-of-the-art (SOTA) results, it may suffer from two drawbacks. First, rather than a static inter-relationship, in real-world applications, the mutual dependencies among entities could be evolving. For instance, in a water treatment plant, when the motorized valve is turned on, it causes the water level to rise. And when the water level exceeds a certain threshold, the motorized valve is closed. These interactions between the actuator of the motorized valve and the water level sensor are strongly coupled. Meanwhile, due to different work conditions, the water level needs to rise for storing water or decrease for draining water. Such relationships are not set in stone, and thus the static graph structure fails to capture this evolving feature. Question 1: Since dependencies among multi-entities are mutual, is there a better choice to model complicated dependencies than the probabilistic model DAG? Second, entities usually have diverse working mechanisms, leading to diverse sparse characteristics when anomalies occur. And, GANF projects all entities into the same distribution. Question 2: Individual anomalous time series from different entities present diverse sparse characteristics. Is it suitable to map all these individual time series to the same latent space (i.e., normal Gaussian distribution)?
@@
In this paper, we propose MTGFlow and MTGFlow_cluster, unsupervised anomaly detection methods for MTS anomaly detection, to answer the above two questions. First, considering the evolving relations among entities, we introduce graph structure learning to model these changeable interdependencies. To learn the dynamic structure, a self-attention module [35] is plugged into our model for its superior performance on quantifying pairwise interaction. Second, we introduce an entity-aware normalizing flow to handle the diverse inherent characteristics among entities. This approach enables us to model entity-specific density estimation, allowing each entity to be associated with a unique target distribution. Consequently, we can estimate the densities of various entities independently. Moreover, the characteristics of entities are strongly connected to their positions in common scenarios. In industrial IoT, sensors within the same production line or workshop have common characteristics due to their exposure to the same industrial processes and operating conditions [36]. For example, temperature and humidity sensors monitor the same area [4, 37]. Similarly, the common wind speed sensors in wind turbines exhibit similar wind speeds and variations in the same regions [38]. In the domain of transportation, traffic flow sensors along the same stretch of highway or road will exhibit similar patterns because they are influenced by similar traffic conditions and weather [39]. Identifying the common information makes for fine-grained representation learning for downstream tasks [40, 41]. Motivated by this, we propose to cluster these entities into different groups. Each group is assigned to the same target distribution while also preserving the distinctiveness among different groups. This results in the creation of MTGFlow_cluster. Such a clustering strategy enables us to efficiently handle entities with similar properties and better model their respective density distributions. Incorporating these proposed technologies provides a powerful framework for tackling the complicated characteristics among diverse entities. However, this fine-grained density estimation of MTGFlow multiplies memory overhead as the number of entities increases. We share entity/cluster-specific model parameters to reduce model size. As a result, MTGFlow and MTGFlow_cluster achieve more fine-grained density estimation without extra memory consumption.
@@ 2.2 Graph Structure Learning for Anomaly Detection
Given a graph structure, graph neural networks  [55, 56] have achieved great success in modeling intrinsic structure patterns. Wang et al. [57] extend the OCSVM to graph anomaly detection, combining the powerful representation ability of GNN. Tang et al. [58] analyze the anomaly detection from the graph spectral domain and find graph anomalies lead to a shift of spectral energy distributions. They generate band-pass filters to capture this energy shift and further perform anomaly detection. However, In real scenarios, the prior graph structure is hard to provide. Therefore, it is important to learn the underlying graph structure via the graph embedding itself [59]. GDN [10] learns a directed graph via node embedding vectors. According to the cosine similarity of embedding vectors, top-K candidates of each node are considered to have dependencies. They rely on either the label supervision or normal training dataset. In the field of unsupervised anomaly detection of MTS, some recent work attempts to explore this area. GANF [32] uses a DAG to model the relationships between multiple sensors and then learns the structure of the DAG through continuous optimization while applying a simplified constraint to facilitate backward propagation Our work, MTGFlow and MTGFlow_cluster, models the mutual complex dependencies as a fully connected graph via a self-attention mechanism, so that a much more flexible relation among entities can be represented.
@@ 2.3 Normalizing Flow for Anomaly Detection
Figure 2: Overview of the proposed MTGFlow and MTGFlow_cluster. Within a sliding window of size $T$, time series $x^{c}$ is fed to the RNN module to capture the temporal correlations. Hidden states of RNN are regarded as time encoding, $H^{c}$. Meanwhile, $x^{c}$ is also input to the graph structure learning module to capture dynamic interdependencies among entities, which are modeled as adjacency matrix $A^{c}$. The spatio-temporal conditions $C^{c}$ are derived via the graph convolution operation for $H^{c}$ and $A^{c}$. Finally, $C^{c}$ is used to help entity/cluster-aware normalizing flow model to produce entity-specific density estimation for the distribution of time series.
@@ 3.1 Self-Attention
$\centering\begin{split}Query=XW^{Query}&\quad\text{and}\quad Key=XW^{Key}\\ Attention\_matrix=&Softmax\left(\frac{QueryKey^{T}}{\sqrt{D}}\right)\end{split}\@add@centering$ (1)
@@ 3.2 Normalizing Flow
Normalizing flow is an unsupervised technique for estimating density, achieved by mapping the original distribution to an arbitrary target distribution using a sequence of invertible affine transformations. When density estimation on original data distribution $\mathcal{X}$ is intractable, an alternative option is to estimate $z$ density on target distribution $\mathcal{Z}$. Specifically, suppose a source sample $x\in\mathcal{R}^{D}\sim\mathcal{X}$ and a target distribution sample $z\in\mathcal{R}^{D}\sim\mathcal{Z}$. Bijective invertible transformation $\mathcal{F}_{\theta}$ aims to achieve one-to-one mapping $z=f_{\theta}(x)$ from $\mathcal{X}$ to $\mathcal{Z}$. According to the change of variable formula, we can get
$P_{\mathcal{X}}(x)=P_{\mathcal{Z}}(z)\left|\det\frac{\partial{f_{\theta}}}{\partial{x}^{T}}\right|$ (2)
Benefiting from the invertibility of mapping functions and tractable jacobian determinants $\left|\det\frac{\partial{f_{\theta}}}{\partial{x}^{T}}\right|$. The objective of flow models is to achieve $\hat{z}=z$, where $\hat{z}=f_{\theta}(x)$. Parameters $\theta$ of ${f}_{\theta}$ can be directly estimated by MLE as follows
@@ 4.2 Overview of MTGFlow and MTGFlow_cluster
The main concepts behind MTGFlow and MTGFlow_cluster are to dynamically model mutual dependencies so that fine-grained density estimation of the multivariate time series can be obtained. This precise estimation allows for better detection of low-density regions, even in datasets with high anomaly contamination during training. The overview of MTGFlow and MTGFlow_cluster can be seen in Fig. 2. Concretely, each entity’s temporal variations are modeled using an RNN, while a graph structure learning module is used to model the dynamic interdependencies. The RNN’s time encoding output is then combined with the graph convolution operation, using the corresponding learned graph structure. These outputs are called spatio-temporal conditions, as they contain both temporal and structural information. In MTGFlow, these conditions are then input into the entity-aware normalizing flow module to achieve precise fine-grained density estimation. And MTGFlow_cluster is designed to first cluster these entities and then estimate the density via cluster-aware normalizing flow. As we do not have the cluster prior, we utilize time series clustering methods like KShape [40] to perform group assignments. Finally, the deviations of $\hat{z}$ and $z$ are measured by log likelihoods, and all modules of MTGFlow are optimized together using MLE.
@@ 4.3 Self-Attention based Graph Structure Learning
$a_{ij}^{c}$ $=\frac{\exp(e_{ij}^{c})}{\sum_{j=1}^{K}exp({e_{ij}^{c}})}\text{,}\quad A^{c}=\begin{bmatrix}a_{11}^{c}&\cdots&a_{1K}^{c}\\ \vdots&\ddots&\vdots\\ a_{K1}^{c}&\cdots&a_{KK}^{c}\end{bmatrix}$ (10)
@@ 4.5 Entity/Cluster-aware Normalizing Flow
$\begin{gathered}P_{\mathcal{X}_{k}}(x_{k})=P_{\mathcal{Z}_{k}}(f_{\theta}^{k}(x_{k}|C))\left|\det\frac{\partial{f_{\theta}^{k}}}{\partial{x_{k}}^{T}}\right|\\ \mathcal{Z}_{k}=\mathcal{N}(\mu_{k},I)\\ \mu_{k}\sim\mathcal{N}(0,1)\\ \end{gathered}$ (12)
@@
$\begin{gathered}P_{\mathcal{X}_{k}}(x_{k})=P_{\mathcal{Z}_{k}}(f_{\theta}^{k}(x_{k}|C))\left|\det\frac{\partial{f_{\theta}^{k}}}{\partial{x_{k}}^{T}}\right|\\ \mathcal{Z}_{k}=\mathcal{N}(\mu_{m},I)\enspace if\enspace k\in cluster_{m}\\ \mu_{m}\sim\mathcal{N}(0,1)\\ {\begin{Bmatrix}cluster_{1},\dots,cluster_{m}\end{Bmatrix}}=KShape\left\{\mathcal{D}\right\}\end{gathered}$ (13)
@@
$P_{\mathcal{X}_{k}}(x_{k})=P_{\mathcal{Z}_{k}}(f_{\theta}(x_{k}|C))\left|\det\frac{\partial{f_{\theta}}}{\partial{x_{k}}^{T}}\right|$ (14)
@@ 5.1.1 Dataset
In addition, we also consider the common OCC setting for a comprehensive evaluation. Under the OCC setting, the training datasets are required to be normal data, so we use the original training datasets of the above-mentioned datasets as training datasets. Besides, Wu et al. [47] provides elaborate and challenging datasets (UCR) including 250-sub univariate datasets from diverse real scenarios (e.g., from insert and NASA spacecraft). For every sub-dataset, the training data is free of anomalies and the test data has only one short-period anomaly. Therefore, we follow the origin split for evaluation on UCR. Although MTGFlow is designed to detect multivariate time series anomaly detection, it can also be applied to univariate series. Note that MTGFlow is equivalent to MTGFlow_cluster when detecting univariate time series.
TABLE II: Anomaly detection performance of AUROC(%) on five public datasets. We test MTGFlow and MTGFlow_cluster in unsupervised and OCC settings.
@@ 5.3.1 Entity-specific Density Estimation
We further explore whether the distributions of all entities are transformed into different target distributions to verify our entity-aware design. Since the window size is 60, the corresponding transformed distributions are also 60-dimensional distributions.
Every single dimension of the multivariate Gaussian distribution is a Gaussian distribution. For better visualization, we present the 0-$th$ dimension of the transformed distributions in Fig. 11.
Nine distributions of different entities are displayed. It can be seen that these distributions have been projected as unique distributions. Moreover, these distributions are successfully converted to preset Gaussian distributions with different mean vectors. The one-to-one mapping models entity-specific distributions and captures their respective sparse characteristics of anomalies, thus mitigating the above-mentioned performance sacrifice.
@@ 6 Conclusion
In this work, we proposed MTGFlow and MTGFlow_cluster, unsupervised anomaly detection approaches for MTS based on the dataset with absolute zero known label. They first model the distribution of training data and then estimate the density of test samples, which are identified anomalies when they lie in low-density regions of this distribution. Extensive experiments on real-world datasets demonstrate its superiority, even if there exists high anomaly contamination. MTGFlow and MTGFlow_cluster achieve SOTA anomaly detection methods for MTS and univariate series in both unsupervised and OCC settings. The superior anomaly detection performance of MTGFlow and MTGFlow_cluster is attributed to dynamic graph structure learning and entity/cluster-aware density estimation. In addition, we explore various interdependencies that exist between individual entities from the learned dynamic graph structure. And a detected anomaly can be understood and localized via entity anomaly scores. In the future, we plan to apply our model to more flow models and further improve the practicality of our approach.

### 07 Feb 2024 v2

@@ 1 Introduction
Density estimation is a promising approach for unsupervised anomaly detection because they do not depend on the assumption that training datasets are all normal. As presented in Fig. 1b, it detects an abnormal sample based on the widely accepted hypothesis that abnormal instances exhibit sparse densities than the normal, i.e., abnormal samples are typically in the low-density regions while normal samples are in the high-density regions [25, 26, 27]. Previous works have attempted this strategy for time series prediction, and the key challenge lies in the accurate density estimation of the distribution. In [28, 29, 30], time series density is modeled as the parameterized probability distribution, while it is still challenging to model a more complex data distribution [31].
@@
More recently, GANF [32] explores the density estimation to tackle MTS anomaly detection task. In their design, the static directed acyclic graph (DAG) is leveraged to model intractable dependence among multiple entities, and normalizing flow is employed to estimate an overall distribution for all entities together [33, 34]. Although GANF has achieved state-of-the-art (SOTA) results, it may suffer from two drawbacks. First, rather than a static inter-relationship, in real-world applications, the mutual dependencies among entities could be evolving. For instance, in a water treatment plant, when the motorized valve is turned on, it causes the water level to rise. And when the water level exceeds a certain threshold, the motorized valve is closed. These interactions between the actuator of the motorized valve and the water level sensor are strongly coupled. Meanwhile, due to different work conditions, the water level needs to rise for storing water or decrease for draining water. Such relationships are not set in stone, and thus the static graph structure fails to capture this evolving feature. Question 1: Since dependencies among multi-entities are mutual, is there a better choice to model complicated dependencies than the probabilistic model DAG? Second, entities usually have diverse working mechanisms, leading to diverse sparse characteristics when anomalies occur. And, GANF projects all entities into the same distribution. Question 2: Individual anomalous time series from different entities present diverse sparse characteristics. Is it suitable to map all these individual time series to the same latent space (i.e., normal Gaussian distribution)?
@@
In this paper, we propose MTGFlow and MTGFlow_cluster, unsupervised anomaly detection methods for MTS anomaly detection, to answer the above two questions. First, considering the evolving relations among entities, we introduce graph structure learning to model these changeable interdependencies. To learn the dynamic structure, a self-attention module [35] is plugged into our model for its superior performance on quantifying pairwise interaction. Second, we introduce an entity-aware normalizing flow to handle the diverse inherent characteristics among entities. This approach enables us to model entity-specific density estimation, allowing each entity to be associated with a unique target distribution. Consequently, we can estimate the densities of various entities independently. Moreover, the characteristics of entities are strongly connected to their positions in common scenarios. In industrial IoT, sensors within the same production line or workshop have common characteristics due to their exposure to the same industrial processes and operating conditions [36]. For example, temperature and humidity sensors monitor the same area [4, 37]. Similarly, the common wind speed sensors in wind turbines exhibit similar wind speeds and variations in the same regions [38]. In the domain of transportation, traffic flow sensors along the same stretch of highway or road will exhibit similar patterns because they are influenced by similar traffic conditions and weather [39]. Identifying the common information makes for fine-grained representation learning for downstream tasks [40, 41]. Motivated by this, we propose to cluster these entities into different groups. Each group is assigned to the same target distribution while also preserving the distinctiveness among different groups. This results in the creation of MTGFlow_cluster. Such a clustering strategy enables us to efficiently handle entities with similar properties and better model their respective density distributions. Incorporating these proposed technologies provides a powerful framework for tackling the complicated characteristics among diverse entities. However, this fine-grained density estimation of MTGFlow and MTGFlow_cluster multiplies memory overhead as the number of entities/clusters increases. We share entity/cluster-specific model parameters to reduce model size. As a result, MTGFlow and MTGFlow_cluster achieve more fine-grained density estimation without extra memory consumption.
@@ 2.2 Graph Structure Learning for Anomaly Detection
Given a graph structure, graph neural networks  [55, 56] have achieved great success in modeling intrinsic structure patterns. Wang et al. [57] extend the OCSVM to graph anomaly detection, combining the powerful representation ability of GNN. Tang et al. [58] analyze the anomaly detection from the graph spectral domain and find graph anomalies lead to a shift of spectral energy distributions. They generate band-pass filters to capture this energy shift and further perform anomaly detection. However, In real scenarios, the prior graph structure is hard to provide. Therefore, it is important to learn the underlying graph structure via the graph embedding itself [59]. GDN [10] learns a directed graph via node embedding vectors. According to the cosine similarity of embedding vectors, top-K candidates of each node are considered to have dependencies. They rely on either the label supervision or normal training dataset. In the field of unsupervised anomaly detection of MTS, some recent work attempts to explore this area. GANF [32] uses a DAG to model the relationships between multiple sensors and then learns the structure of the DAG through continuous optimization while applying a simplified constraint to facilitate backward propagation. Our work, namely MTGFlow and MTGFlow_cluster, models the mutual complex dependencies as a fully connected dynamic graph via a self-attention mechanism, so that a much more flexible relation among entities can be represented.
@@ 2.3 Normalizing Flow for Anomaly Detection
Figure 2: Overview of the proposed MTGFlow and MTGFlow_cluster. Within a sliding window of size $T$, time series $x^{c}$ is fed to the RNN module to capture the temporal correlations. Hidden states of RNN are regarded as time encoding, $H^{c}$. Meanwhile, $x^{c}$ is also input to the graph structure learning module to capture dynamic interdependencies among entities, which are modeled as adjacency matrix $A^{c}$. The spatio-temporal conditions $C^{c}$ are derived via the graph convolution operation for $H^{c}$ and $A^{c}$. Finally, $C^{c}$ is used to help entity/cluster-aware normalizing flow model to produce entity/cluster-specific density estimation for the distribution of time series.
@@ 3.1 Self-Attention
$\centering\begin{split}Query=XW^{Query}&\quad\text{and}\quad Key=XW^{Key}\\ Attention\_matrix=&Softmax\left(\frac{QueryKey^{T}}{\sqrt{D}}\right).\end{split}\@add@centering$ (1)
@@ 3.2 Normalizing Flow
Normalizing flow is an unsupervised technique for estimating density, achieved by mapping the original distribution to an arbitrary target distribution using a sequence of invertible affine transformations. When density estimation on original data distribution $\mathcal{X}$ is intractable, an alternative option is to estimate $z$ density on target distribution $\mathcal{Z}$. Let us define a source sample $x\in\mathcal{R}^{D}\sim\mathcal{X}$ and a target distribution sample $z\in\mathcal{R}^{D}\sim\mathcal{Z}$. Bijective invertible transformation $\mathcal{F}_{\theta}$ aims to achieve one-to-one mapping $z=f_{\theta}(x)$ from $\mathcal{X}$ to $\mathcal{Z}$. According to the change of variable formula, we can get
$P_{\mathcal{X}}(x)=P_{\mathcal{Z}}(z)\left|\det\frac{\partial{f_{\theta}}}{\partial{x}^{T}}\right|.$ (2)
Benefiting from tractable jacobian determinants $\left|\det\frac{\partial{f_{\theta}}}{\partial{x}^{T}}\right|$, the objective of flow models is to achieve $\hat{z}=z$, where $\hat{z}=f_{\theta}(x)$. Parameters $\theta$ of ${f}_{\theta}$ can be directly estimated by MLE as follows
@@ 4.2 Overview of MTGFlow and MTGFlow_cluster
The main concepts behind MTGFlow and MTGFlow_cluster are to dynamically model mutual dependencies so that fine-grained density estimation of the multivariate time series can be obtained. This precise estimation allows for better detection of low-density regions, even in datasets with high anomaly contamination during training. The overview of MTGFlow and MTGFlow_cluster can be seen in Fig. 2. Concretely, each entity’s temporal variations are modeled using an RNN, while a graph structure learning module is used to model the dynamic interdependencies. The RNN’s time encoding output is then combined with the graph convolution operation, using the corresponding learned graph structure. These outputs are called spatio-temporal conditions, as they contain both temporal and structural information. In MTGFlow, these conditions are then input into the entity-aware normalizing flow module to achieve precise fine-grained density estimation. And MTGFlow_cluster is designed to first cluster these entities and then estimate the density via cluster-aware normalizing flow. As we do not have the cluster prior, we utilize time series clustering methods like KShape [40] to perform group assignments. Finally, the deviations of $\hat{z}$ and $z$ are measured by log likelihoods, and all modules of MTGFlow and MTGFlow_cluster are optimized together using MLE.
@@ 4.3 Self-Attention based Graph Structure Learning
$a_{ij}^{c}$ $=\frac{\exp(e_{ij}^{c})}{\sum_{j=1}^{K}exp({e_{ij}^{c}})}\text{,}\quad A^{c}=\begin{bmatrix}a_{11}^{c}&\cdots&a_{1K}^{c}\\ \vdots&\ddots&\vdots\\ a_{K1}^{c}&\cdots&a_{KK}^{c}\end{bmatrix}.$ (10)
@@ 4.5 Entity/Cluster-aware Normalizing Flow
$\begin{gathered}P_{\mathcal{X}_{k}}(x_{k})=P_{\mathcal{Z}_{k}}(f_{\theta}^{k}(x_{k}|C))\left|\det\frac{\partial{f_{\theta}^{k}}}{\partial{x_{k}}^{T}}\right|,\\ \mathcal{Z}_{k}=\mathcal{N}(\mu_{k},I),\\ \mu_{k}\sim\mathcal{N}(0,1).\\ \end{gathered}$ (12)
@@
$\begin{gathered}P_{\mathcal{X}_{k}}(x_{k})=P_{\mathcal{Z}_{k}}(f_{\theta}^{k}(x_{k}|C))\left|\det\frac{\partial{f_{\theta}^{k}}}{\partial{x_{k}}^{T}}\right|,\\ \mathcal{Z}_{k}=\mathcal{N}(\mu_{m},I)\enspace if\enspace k\in cluster_{m},\\ \mu_{m}\sim\mathcal{N}(0,1),\\ {\begin{Bmatrix}cluster_{1},\dots,cluster_{m}\end{Bmatrix}}=KShape\left\{\mathcal{D}\right\}.\end{gathered}$ (13)
@@
$P_{\mathcal{X}_{k}}(x_{k})=P_{\mathcal{Z}_{k}}(f_{\theta}(x_{k}|C))\left|\det\frac{\partial{f_{\theta}}}{\partial{x_{k}}^{T}}\right|.$ (14)
@@ 5.1.1 Dataset
In addition, we also consider the common OCC setting for a comprehensive evaluation. Under the OCC setting, the training datasets are required to be normal data, so we use the original training datasets of the above-mentioned datasets as training datasets. Besides, Wu et al. [47] provides elaborate and challenging datasets (UCR) including 250-sub univariate datasets from diverse real scenarios (e.g., from biology and industry). For every sub-dataset, the training data is free of anomalies and the test data has only one short-period anomaly. Therefore, we follow the origin split for evaluation on UCR. Although MTGFlow is designed to detect multivariate time series anomaly detection, it can also be applied to univariate series. Note that MTGFlow is equivalent to MTGFlow_cluster when detecting univariate time series.
TABLE II: Anomaly detection performance of AUROC(%) on five public datasets. We test MTGFlow and MTGFlow_cluster in unsupervised and OCC settings. The best results are highlighted in bold.
@@ 5.3.1 Entity-specific Density Estimation
We further explore whether the distributions of all entities are transformed into different target distributions to verify our entity-aware design. Since the window size is 60, the corresponding transformed distributions are also 60-dimensional distributions. Every single dimension of the multivariate Gaussian distribution is a Gaussian distribution. For better visualization, we present the 0-$th$ dimension of the transformed distributions across the whole test dataset in Fig. 11. Nine distributions of different entities are displayed. It can be seen that these distributions have been projected as unique distributions. Moreover, these distributions are successfully converted to preset Gaussian distributions with different mean vectors. The one-to-one mapping models entity-specific distributions and captures their respective sparse characteristics of anomalies, thus mitigating the above-mentioned performance sacrifice.
@@ 6 Conclusion
In this work, we proposed MTGFlow and MTGFlow_cluster, unsupervised anomaly detection approaches for MTS based on the dataset with absolute zero known label. They first model the distribution of training data and then estimate the density of test samples, which are identified anomalies when they lie in low-density regions of this distribution. Extensive experiments on real-world datasets demonstrate their superiority, even if there exists high anomaly contamination. MTGFlow and MTGFlow_cluster achieve SOTA anomaly detection methods for MTS and univariate series in both unsupervised and OCC settings. The superior anomaly detection performance of MTGFlow and MTGFlow_cluster is attributed to dynamic graph structure learning and entity/cluster-aware density estimation. In addition, we explore various interdependencies that exist between individual entities from the learned dynamic graph structure. And a detected anomaly can be understood and localized via entity anomaly scores. In the future, we plan to apply our model to more flow models and further improve the practicality of our approach.